thatkid101:

C^2+40c+400 , factor the perfect square trinomial

SmokeyBrown:

This is a nice example problem, because it's simple to see that 20+20 = 40 and that 20*20 = 400. Based on this insight, we can reason almost intuitively that the perfect square of c^2+40c+400 would be (c + 20) We can even check our work by multiplying (c+20) * (c+20) in order to get the result c^2 + 40c + 400

thatkid101:

so c+20^2

SmokeyBrown:

In a more general case, if there exists a perfect square for a polynomial in the form ax^2 + bx + c, the perfect square will take the form (dx + e) such that d*d = a; de + de = b; and e*e = c. In this case, the perfect square of c^2 + 40c + 400 is (c + 20) because 1*1 = 1 1*20 + 1*20 = 40 and 20 * 20 = 400 You could also check the work by performing (c+20)^2 or (c+20) * (c+20), of course

SmokeyBrown:

@thatkid101 wrote:
so c+20^2
And it is important to distinguish (c+20)^2 rather than c+20^2; The first will get you the result you want, c^2 + 40c + 400; The second will get you c+400, since that notation only squares the 20 by itself

thatkid101:

this hurts my brain ;-;

SmokeyBrown:

@thatkid101 wrote:
this hurts my brain ;-;
I may have used too many different letters in my explanation. That certainly didn't make it easy to keep track of what I tried to communicate. We can work through it in a more straightforward way by simply performing the multiplication: (c+20) * (c+20) This can also be rewritten as c*(c+20) + 20*(c+20), according to the distributive property, if that makes it easier to visualize. The result is: c^2 + 20c + 20c + 400, which resolves to c^2 + 40c +400 So, we've shown that we can multiply (c+20) by itself in order to get our original polynomial. By definition, that means (c+20) is a perfect square of the original polynomial (or maybe vice-versa?) In any case, I hope this helps demonstrate the point I was trying to make. Please let me know if you have any questions :)

thatkid101:

could u help me with some other questions ?

Frostsoldier101:

ok

Frostsoldier101:

wait who

thatkid101:

either or idc who i just need help

Frostsoldier101:

ok me

thatkid101:

factor the perfect square trinomial. 9m^2-42m+49

Frostsoldier101:

(a+b)2=a2+2ab+b2.

SmokeyBrown:

@frostsoldier101 wrote:
(a+b)2=a2+2ab+b2.
A much better way to communicate what I tried to in my own convoluted way up top :)

Frostsoldier101:

lol

Frostsoldier101:

ight bye

thatkid101:

thank u ^^

Frostsoldier101:

you welcome